Journalist Report – May 13th

Today was Sol 8.

Today started with a relaxed morning, the crew members worked on their own projects. There was the continuation of work on processing the photographs, more progress was made to recover the software of the rover with another 2am conference with the UK, and another batch of bacteria was mixed with soil to re-inoculate the seeds, the previous equipment has been sterilised and the bacteria has been killed.

An EVA was taken today which was 2 ½ hours long, it was an eventful EVA that was mixed with an engineering EVA. Firstly, we delivered another Mars mouse to the release point, after we went to the experiment site and started the two experiments. The first was to test the oxygen levels in the suits; a circuit was completed to show the differences in the oxygen levels. However, due to the air holes and lack of sealing a new experiment must be proposed. The drone experiment continued, with the testing the X-5 and the Palantir. A failure occurred in the Palantir during piloting, the damage seems unrepairable. Unfortunately, the EVA finished with a failure of Opportunity. Opportunity had to be towed back to the Hab during and engineering EVA.

Copyright © The Mars Society. All rights reserved. | Main Site